OpenStudy (anonymous):

Find the slope of this line 5x+3=3y

OpenStudy (hoblos):

3y = 5x+3 y=(5/3)x +1 slope = 5/3

OpenStudy (anonymous):

Thanks so much

OpenStudy (hoblos):

my pleasure :D

OpenStudy (anonymous):

sent you a message back